OpenStudy (anonymous):

1. Solve the system by substitution 2x+y=-11 3x-4y=11 A. (3,5) B. (-5,-3) C. (-3,-5) D. (5,3)

OpenStudy (azureilai):

2x+y=-11 y=-2x-11 3x-4(-2x-11)=11 Now, solve the above equation

OpenStudy (anonymous):

Look, we have : 2x+y=-11 3x-4y=11 ________________________ +4(2x+y=-11) 3x-4y=11 ________________ have : 8x+4y=-44 3x-4y=11 Sum them : 8x+4y+3x-4y -44+11 have : 11x=-33 and x=3 To find y you can get 3 to x in one of the equations and find that :)

OpenStudy (anonymous):

3. What is the solution of the following system? -3x-2y=-12 9x+6y=-9 A. (2, 1) B. no solutions C. (–2, –1) D. infinitely many solutions

OpenStudy (anonymous):

Thanks